# LIFT: A Novel Framework for Enhancing Long-Context Understanding of LLMs via Long Input Fine-Tuning

### v4

@@ 2 Related Work
Retrieval-Augmented Generation (RAG). RAG (Lewis et al., 2020) improves the performance of long-context understanding by integrating LLMs with external data sources for retrieval (Xu et al., 2023; Jiang et al., 2024c; Wang et al., 2024a; Jin et al., 2024; Yue et al., 2025; Edge et al., 2024), thereby avoiding the need to feed the entire long input. Its performance heavily relies on the quality of retrieved content, which must be relevant and concise enough to fit within models’ short context windows. RAG can experience significant performance degradation or hallucination issues when the retrieved context is inaccurate or mismatched. A comparison of our LIFT with RAG and long-context ICL is in Table 1.
@@ 4 Experiments
Table 3: Performance of different methods on LooGLE based on the Llama-3-8B-Instruct. We evaluate the accuracy of the methods on LooGLE short-dependency QA (ShortQA) and long-dependency QA (LongQA). Comprehension & reasoning, multiple info retrieval, computation, and timeline reorder are the subtasks in LongQA and we evaluate the accuracy on each of them.
MemoryLLM 33.06 20.44 29.31 15.53 8.00 18.14
LlamaIndex 41.93 21.07 33.00 17.11 12.00 9.77
TempLoRA 29.67 25.97 32.02 19.47 12.00 32.56
LLMLingua 24.14 23.80 33.50 20.00 8.00 19.53
Truncated ICL 44.49 15.44 25.37 15.26 5.00 1.86
Finetune-Raw 33.93 25.16 30.05 20.53 18.00 27.44
LIFT with 5QA 45.67 26.79 29.80 21.58 14.00 36.28
LIFT with 10QA 52.69 27.25 27.83 22.63 16.00 39.53
@@ Appendix D Comparison against LLoCO
Table 7: Comparison between LIFT and LLoCO on QuALITY.

### v5

@@ 1 Introduction
  • Flexibility in task adaptation. LIFT is a versatile framework that imposes no restrictions on the synthetic task generation strategy, allowing it to adapt to diverse downstream requirements. Beyond long-context comprehension, LIFT is applicable to a wide range of scenarios, such as in-context skill acquisition (Appendix K), allowing LLMs to understand and learn from contexts.
@@ 2 Related Work
Retrieval-Augmented Generation (RAG). RAG (Lewis et al., 2020) improves the performance of long-context understanding by integrating LLMs with external data sources for retrieval (Xu et al., 2023; Jiang et al., 2024c; Wang et al., 2024a; Jin et al., 2024; Yue et al., 2025; Edge et al., 2024), thereby avoiding the need to feed the entire long input. Its performance heavily relies on the quality of retrieved content, which must be relevant and concise enough to fit within models’ short context windows. RAG can experience significant performance degradation or hallucination issues when the retrieved context is inaccurate or mismatched. This issue becomes particularly severe when downstream queries differ substantially from the document’s wording. Such situations are common in practice, including enterprise compliance review, where queries describe detailed scenarios while the governing rules are abstract and conclusive, and complex logical reasoning, where the context consists of the model’s exploratory trials. In these cases, training can internalize abstract heuristics, whereas such learned heuristics are difficult for RAG to retrieve. A comparison of our LIFT with RAG and long-context ICL is in Table 1.
@@ 4 Experiments
Table 3: Performance of different methods on LooGLE based on the Llama-3-8B-Instruct. We evaluate the accuracy of the methods on LooGLE short-dependency QA (ShortQA) and long-dependency QA (LongQA). Timeline reorder, multiple info retrieval, computation, and comprehension & reasoning are the subtasks in LongQA and we evaluate the accuracy on each of them.
Methods ShortQA LongQA Timeline reorder Multiple info retrieval Computation Comprehension & Reasoning
MemoryLLM 33.06 20.44 18.14 15.53 8.00 29.31
LlamaIndex 41.93 21.07 9.77 17.11 12.00 33.00
TempLoRA 29.67 25.97 32.56 19.47 12.00 32.02
LLMLingua 24.14 23.80 19.53 20.00 8.00 33.50
Truncated ICL 44.49 15.44 1.86 15.26 5.00 25.37
Finetune-Raw 33.93 25.16 27.44 20.53 18.00 30.05
LIFT with 5QA 45.67 26.79 36.28 21.58 14.00 29.80
LIFT with 10QA 52.69 27.25 39.53 22.63 16.00 27.83
@@ Appendix D Comparison against LLoCO
Table 7: Comparison between LIFT and LLoCO on LooGLE.
@@ Appendix H Additional Experiments on the Variants of LIFT
Appendix I Analysis of the Quality of Synthetic Data
Since LIFT relies on synthetic data to memorize and understand context, it’s necessary to inspect the quality of the generated synthetic data. We examine both the format and content diversity of the synthetic tasks generated by Qwen2.5-72B for LooGLE benchmark. To assess format diversity, we prompt Qwen2.5-72B to rewrite all QA pairs into “what-is” style; to assess content diversity, we categorize the synthetic tasks into “event”, “location”, “person”, “time”, and “other” using Qwen2.5-72B and fine-tune the model on each category separately. The proportions of the categories are shown in Figure 7 and the evaluation results are illustrated in Table 11. The results indicate that diversity matters, but the generator is capable of producing diverse QA pairs.
Figure 7: Distribution of Synthetic QA Tasks by Content Category.
Table 11: Performance on LooGLE of Llama-3-8B-Instruct fine-tuned on synthetic tasks formatted as “what-is” questions, or on synthetic tasks from only one of the five task categories.
What-is Event Location Person Time Other
ShortQA 40.11 33.84 25.70 40.54 23.34 23.34
LongQA 13.78 17.82 16.37 17.91 21.44 13.93
The following are the examples of each category:
Event
Q
: What unique item did Picardo design for the fashion brand Loewe in 1959?
A: He designed a pack of playing cards (baraja de naipes) for their exclusive advertising use. Location
Q
: Where was José Luis Picardo born?
A: He was born in Jerez de la Frontera, located in the Province of Cádiz, Andalucía, Spain. Person
Q
: Whose architectural studio did Picardo join to avoid evacuation during the Spanish Civil War?
A: He joined the studio of architect and professor Luis Moya Blanco. Time
Q
: When did the Parador de Guadalupe officially open to the public?
A: The hotel opened on December 11, 1965. Other
Q
: What was the significance of the ”Manifiesto de la Alhambra” that Picardo signed in 1952?
A: It is considered one of the most remarkable texts in 20th-century Spanish architectural historiography and sought inspiration from the Alhambra for a distinctively Spanish form of modern architecture.
Appendix J Using the Target Model as Synthetic Data Generator
In our paper, we use Qwen2.5-72B as the generator, which is capable of producing diverse and high-quality QA pairs. However, a large generator is not necessary to guarantee the quality of the synthetic data. Moreover, LIFT’s advantage stems from its understanding of context, rather than from knowledge distilled from a strong generator. To verify this, we conducted an additional experiment in which the target model itself (Meta-Llama-3-8B-Instruct) serves as the generator. The results are shown in Table 12, which suggest that an 8B model is already sufficient to generate useful synthetic tasks, and that the effectiveness of LIFT does not critically depend on a strong generator.
Table 12: Comparison of LIFT with the target model (Llama-3-8B-Instruct) itself as the generator and the ICL baseline.
ICL 44.49 15.44 25.37 15.26 5.00 1.86
LIFT 48.49 26.93 34.14 21.49 15.00 28.32
Appendix K Extending LIFT to Diverse Task Formats
To evaluate the generalizability of LIFT beyond document-based question answering (e.g., SQuAD and LooGLE), we assess its performance on document summarization (GovReport, Huang et al. (2021)) and a specialized in-context skill acquisition scenario derived from GSM8K. This setup tests whether LIFT can internalize complex reasoning patterns and structural constraints into model parameters, rather than merely retrieving facts.
In the GSM8K skill acquisition task, the input context consists of 10 solved examples. The model must learn the underlying reasoning logic from these examples to solve a separate, unseen test case. For both benchmarks, we compare LIFT against a truncated in-context learning baseline. We utilize Qwen2.5-72B-Instruct as the synthetic task generator and Llama-3-8B-Instruct as the target model.
For LIFT, we design task-specific strategies to generate synthetic tasks:
  • GovReport: The generator is prompted to produce 3 summarizations of their granularity from coarse to fine every 20 sentences. The target model is fine-tuned on the summarizations.
  • GSM8K: For each of the 10 examples, the generator extracts the reasoning steps and rewrites them while keeping the meaning. Moreover, it synthesizes 10 more examples sharing a similar solving strategy with the example.
As shown in Table 13, LIFT consistently outperforms the ICL baseline. Notably, the significant gain on GSM8K (+7.8%) demonstrates that LIFT’s synthetic task generation strategy is superior to in-context learning, allowing the model to learn from contexts better.
Table 13: Performance comparison between LIFT and the truncated ICL baseline on GovReport and the in-context skill-acquisition task (GSM8K).
Method GovReport GSM8K
ICL 29.1 57.6
LIFT 30.9 65.4